Find the  ∑ (x − μ)2
 on the values of (3, 5, 8,9) with the sample size of n=3.





1
Expert's answer
2021-11-12T06:50:50-0500


We have population values "3,5,8,9," population size "N=4"

"\mu=\dfrac{3+5+8+9}{4}=6.25""\sigma^2=\dfrac{1}{4}((3-6.25)^2+(5-6.25)^2+(8-6.25)^2"

"+(9-6.25)^2)=5.6875"

We have population values "3,5,8,19," population size "N=4" and sample size "n=3." Thus, the number of possible samples which can be drawn without replacement is



"\dbinom{N}{n}=\dbinom{4}{3}=4""\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 3,5,8 & 16/3 \\
 \hdashline
 2 & 3,5,9 & 17/3 \\
 \hdashline
 3 & 3,8,9 & 20/3 \\
 \hdashline
 4 & 5,8,9 & 22/3 \\
 \hline
\end{array}"

The sampling distribution of the sample means.



"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
& \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 & 16/3 & 1 & 1/4 & 6.25 & 256/36 \\
 \hdashline
 & 17/3 & 1& 1/4 & 17/12 & 289/36 \\
 \hdashline
 & 20/3 & 1 & 1/4 & 20/12 & 400/36 \\
 \hdashline
 & 22/3 & 1& 1/4 & 22/12 & 484/36 \\
 \hdashline
 Total & & 4 & 1 & 6.25 & 1429/36 \\ \hline
\end{array}"



The mean of the sampling distribution of the sample means is equal to the the mean of the population.


"E(\bar{X})=\mu_{\bar{X}}=6.25=\mu"




"Var(\bar{X})=\sum(\bar{X}-\mu)^2"




"=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"


"=\dfrac{1429}{36}-(6.25)^2=\dfrac{91}{144}"



Verification:


"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{5.6875}{3}(\dfrac{4-3}{4-1})"




"=\dfrac{91}{144}, True"




"\sum(\bar{X}-\mu)^2=\dfrac{91}{144}"